OpenStudy (anonymous):

Can anyone explain to me how to find vertical asmptotes and holes for any graph the one i have is like (X+1)(X-2)*OVER*(X-2)(X+4)

jimthompson5910 (jim_thompson5910):

Does this simplify at all?

OpenStudy (anonymous):

uh i have no idea

jimthompson5910 (jim_thompson5910):

do you see any common factors

OpenStudy (anonymous):

Only X

jimthompson5910 (jim_thompson5910):

how about x-2?

jimthompson5910 (jim_thompson5910):

that appears twice

OpenStudy (anonymous):

Oh So..this might be much to ask but can you explain how to solve it?...Sorry :/

jimthompson5910 (jim_thompson5910):

\[\Large \frac{(x+1)(x-2)}{(x-2)(x+4)}\] \[\Large \frac{(x+1)\cancel{(x-2)}}{\cancel{(x-2)}(x+4)}\] \[\Large \frac{x+1}{x+4}\]

jimthompson5910 (jim_thompson5910):

See how I'm cancelling those common factors?

OpenStudy (anonymous):

I do

jimthompson5910 (jim_thompson5910):

So \[\Large \frac{(x+1)(x-2)}{(x-2)(x+4)}\] simplifies to \[\Large \frac{x+1}{x+4}\]

jimthompson5910 (jim_thompson5910):

In other words, they are the same

jimthompson5910 (jim_thompson5910):

But, in order for them to be the exact same, the domains must match So we must specify that \(\Large x \neq 2\) and \(\Large x \neq -4\) to avoid dividing by zero

jimthompson5910 (jim_thompson5910):

The vertical asymptote will be x = -4 because this causes a division by zero error in \[\Large \frac{x+1}{x+4}\] and there is a hole at x = 2 because there is no division by zero error in \[\Large \frac{x+1}{x+4}\] but we must make this exclusion to make sure the domains match

OpenStudy (anonymous):

is +2 considered the hole and -4 the asmptote?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Thank you Very much! Im understanding alot better now! i might be asking for a bit more help soon in the future! Please keep an Eye out! Thanks again!

jimthompson5910 (jim_thompson5910):

alright, yw